OpenStudy (anonymous):

find derivative of (2/t^2)+(1/t)+6-12t+20t

OpenStudy (anonymous):

\[f(t) = \frac{ 2 }{ t ^{2} }+\frac{1}{t} +6 - 12t +20t\] first, you'll want to simplify your equation by combinging the like terms. \[f(t) = \frac{ 2 }{ t ^{2} }+\frac{1}{t} +6 +8t\] Then you can format it in a way that'll make it easy to use the chain rule. \[f(t) = 2 { t ^{-2} }+t^{-1} +6 +8t\] now you can find f'(t) by differentiating through. Here's a general example of what you want to do f(t) = a*t^{b} f'(t) = a*(b-1)t^{b-1} Do this for everything in the equation. If it helps you visualize, \[f(t) = 2 { t ^{-2} }+t^{-1} +6 +8t\] is equivalent to \[f(t) = 2 { t ^{-2} }+t^{-1} +6t^{0} +8t\]

OpenStudy (anonymous):

thanks so much, unfortunatly i miss typed the problem because it's 20t^2 instead of just 20t...